### v1

# TurBLiMP: A Turkish Benchmark of Linguistic Minimal Pairs

Abstract
We introduce TurBLiMP, the first Turkish benchmark of linguistic minimal pairs, designed to evaluate the linguistic abilities of monolingual and multilingual language models (LMs). Covering 16 linguistic phenomena with 1000 minimal pairs each, TurBLiMP fills an important gap in linguistic evaluation resources for Turkish. In designing the benchmark, we give extra attention to two properties of Turkish that remain understudied in current syntactic evaluations of LMs, namely word order flexibility and subordination through morphological processes. Our experiments on a wide range of LMs and a newly collected set of human acceptability judgments reveal that even cutting-edge Large LMs still struggle with grammatical phenomena that are not challenging for humans, and may also exhibit different sensitivities to word order and morphological complexity compared to humans.
[…]
TurBLiMP: A Turkish Benchmark of Linguistic Minimal Pairs
Ezgi Başar  Francesca Padovani  Jaap Jumelet  Arianna Bisazza Center for Language and Cognition (CLCG), University of Groningen e.basar@student.rug.nl {f.padovani, j.w.d.jumelet, a.bisazza}@rug.nl
1 Introduction
A foundational insight in linguistics research is that applying minimal changes to a sentence can render it entirely acceptable or unacceptable to native speakers (Chomsky, 1965). Minimal pairs, as illustrated in Example (1), are a widely used diagnostic tool in linguistics.
[…]
Minimal pairs have been a cornerstone of linguistic analysis for decades, and in recent years they have become a vital tool for the linguistic evaluation of language models (LMs). Warstadt et al. (2020) published the first large-scale English Benchmark of Linguistic Minimal Pairs (BLiMP) in an effort to systematically evaluate the linguistic knowledge of language models, and since then various benchmarks have been introduced for other languages.
We contribute to this growing collection by introducing the first Turkish benchmark of linguistic minimal pairs. TurBLiMP enriches the typological diversity of available linguistic evaluation benchmarks by incorporating a morphologically rich agglutinative language with highly flexible word order. While Turkish and other agglutinative languages like Finnish have been the object of several studies focusing on word-level morphology (Ismayilzada et al., 2025), the effects of word order flexibility and morphological complexity on the robustness of sentence-level grammatical judgments have not been studied in detail before. We fill this gap by introducing two sets of experimental minimal pair paradigms.
Our evaluation shows that even top-performing LMs suffer performance losses under word order or subordination manipulations, revealing sensitivities that would otherwise go undetected. Compared to the acceptability judgments we collected from native speakers, baseline tests across 13 models and 16 Turkish phenomena demonstrate that Large LMs can struggle with linguistic tasks where humans perform reliably. By providing this resource, we aim to facilitate linguistically motivated NLP research and contribute a high-quality dataset for linguists and NLP researchers.
2 Minimal Pair Benchmarks
Minimal pairs have played an important role for evaluating the linguistic abilities of language models, targeting phenomena such as subject-verb agreement (Linzen et al., 2016), filler-gap dependencies (Wilcox et al., 2018), and negative polarity items (Jumelet and Hupkes, 2018). Warstadt et al. (2020) then established an English benchmark of 67,000 sentence pairs testing 67 paradigms through automated generation based on linguist-curated templates. This work inspired numerous adaptations for other languages, each employing different benchmark creation strategies. Benchmarks using a similar template-based approach as BLiMP include CLiMP (Chinese, Xiang et al., 2021), ZhoBLiMP (Chinese, Liu et al., 2024), BLiMP-NL (Dutch, Suijkerbuijk et al., 2025), and for Basque/Swahili/Hindi by Kryvosheieva and Levy (2025). Another approach is based on modifying Universal Dependency trees, which has been used by SLING (Chinese, Song et al., 2022), RuBLiMP (Russian, Taktasheva et al., 2024), and MultiBLiMP (Jumelet et al., 2025), a multilingual benchmark covering 101 languages. Other approaches include the extraction of minimal pairs from linguistics journals, employed by JBLiMP (Japanese, Someya and Oseki, 2023), manual creation of pairs, as done for Icelandic by Ármannsson et al. (2025), and the usage of LLMs for generating pairs, as done for Tamil and Indonesian by Leong et al. (2023).
Methodological innovations across these benchmarks reveal key trade-offs between scale, linguistic coverage, and data quality. Template-based generation enables large datasets but risks producing unnatural sentences (Vázquez Martínez et al., 2023), while manual extraction from literature or learner corpora ensures quality at the cost of scale. Some of the benchmarks incorporate hybrid approaches and human validation steps to balance these concerns. TurBLiMP too is the result of such hybrid approaches. While creating our benchmark, we developed strategies specifically adapted to the challenges of creating minimal pairs for Turkish.
3 Turkish Morphosyntax & NLP
Turkish presents a particularly interesting case for BLiMP-style evaluation due to its flexible word order and rich morphological system. Turkish syntactically licenses all six possible orderings of the main sentence constituents: Subject-Object-Verb (SOV) represents the canonical order, while other permutations introduce subtle pragmatic variations without altering the core meaning of the sentence. As a result, evaluating LMs on a language like Turkish makes it possible to test them for their robustness to different positional patterns or grammatical hierarchies, in a way that is not possible with English and other fixed-order languages that dominate the training material of current LLMs.
Furthermore, Turkish has highly productive agglutinative morphology, whereby words typically consist of several morphemes attached to a root. Speakers can easily produce and understand numerous legitimate but low-frequency word forms through regular morphological processes, yielding substantially larger vocabulary requirements for LMs compared to analytic and fusional languages. Many syntactic phenomena are realized in Turkish through morphology, rather than by separate function words like in English and other Indo-European languages that form a large chunk of the world’s highest-resource languages. A salient example is subordination, which largely involves the use of suffixes to nominalize or adverbialize the verb of the embedded clause. For instance, the sentence I know that Elif likes Gaye translates to Elif’in Gaye’yi sevdiğini biliyorum whose structure can be intuitively conveyed as ‘I know the liking of Gaye by Elif’. Here, the nominalized verb ‘like’ takes an accusative case suffix as the object of ‘know’, but also a possessive agreement suffix corresponding to the genitive suffix taken by the subordinate subject ‘Elif’.
In general, agglutinative languages have been shown to be particularly challenging for neural models Gerz et al. (2018); Cotterell et al. (2016); Park et al. (2021); Arnett and Bergen (2025). Focusing on Turkish, Ataman et al. (2017) established that fixed vocabulary constraints combined with suboptimal sub-word segmentation significantly impair neural machine translation performance for agglutinative languages. Ismayilzada et al. (2025) studied LLMs’ ability to produce and systematically understand novel well-formed combinations of morphemes in Turkish and Finnish, and reported limited morphological generalization. These findings suggest that studying flexible-order, morphologically rich languages like Turkish can provide unique insights into the true linguistic capabilities of LMs beyond surface fluency.
4 TurBLiMP
The creation of the TurBLiMP benchmark was motivated by the need for a controlled evaluation benchmark that accounts for the unique linguistic properties of Turkish. Some of these properties include flexible word order, morphological richness, optional pro-drop, and syncretism in third-person subject-verb agreement markers. We now provide a brief linguistic background on our minimal pairs.
Phenomenon Minimal pair Translation
Anaphor Agreement Gezi rota-sın-ı [kendi-miz /*kendi-niz] internet-e bak-ma-dan oluştur-du-k. trip route-3sg.poss-acc [self-1pl.poss /*self-2pl.poss] internet-dat look-neg-abl create-pst-1pl We created the trip itinerary [ourselves / *yourselves] without checking the internet.
Arg. Struct. Trans. Eş-im-in [zevk-in-e /*zevk-in-i] çok güven-ir-im. spouse-1sg.poss-3sg.gen [taste-3sg.poss-dat /*taste-3sg.poss-acc] very trust-AOR-1sg I trust my wife’s taste a lot.
Arg. Struct. Ditrans. Öğretmen [öğrenci-ler-e /*öğrenci-ler-i] yeni konu-yu anlat-tı. teacher [student-pl-dat /*student-pl-acc] new subject-acc explain-pst The teacher explained the new topic to the students.
Binding Yaz tatil-in-de [kendi-m-i /*ben-i] rahatlamış hissed-iyor-um. summer holiday-3.poss-loc [self-1sg-acc /*me] relaxed feel-prog-1sg I feel relaxed during the summer holidays.
Determiners Geçen hafta tad-ı damağ-ım-da kal-an [bir /$\varnothing$] tatlı ye-di-m. last week taste-acc palate-1sg.poss-loc stay-part [a /$\varnothing$] dessert eat-pst-1sg Last week, I ate a dessert with a taste that lingered on my tongue.
Ellipsis Mağaza-da ceket-i Pelin ve [pantolonu Cem /*Cem pantolonu] seç-ti. store-loc jacket-acc Pelin and trouser-acc Cem choose-pst In the store, Pelin chose the jacket and Cem chose the pants.
Irregular Forms Güneş gör-me-yen petunya-lar hemen [ölür/*öler]. sun see-neg-part petunia-pl immediately dies Petunias that do not see the sun die immediately.
Island Effects [Neyi /*Onu neden] dükkan-a getir-en eleman azar işit-ti? [what /*it why] shop-dat bring-part worker scolding hear-pst The worker who brought what to the store was scolded?
Nominalization Konu-nun tekrar [tartış-ıl-ma-sın-ı /*tartış-ıl-dığ-ın-ı] öner-iyor-um. matter-gen again [discuss-pass-MA-poss-acc /*discuss-pass-DIK-poss-acc] suggest-prog-1sg I suggest that the matter be discussed again.
NPI Licensing Kalabalığ-ın ön-ün-de [$\varnothing$ /*hiç] şarkı söyle-di-m. crowd-gen front-poss-loc [$\varnothing$ /*ever] song sing-pst-1sg I (*ever) sang in front of a crowd.
[…]
Quantifiers Mağaza-da [$\varnothing$ /*çoğu] ayakkabı dene-di-m. store-loc [$\varnothing$ /*çoğu] shoe try_on-pst-1sg I tried on shoes in the store.
Relative Clauses Sınav-da [gözetmen-in /*gözetmen-i] uyar-dığ-ı öğrenci yer-in-e geç-ti. exam-loc [proctor-3sg.gen /*proctor-acc] warn-part-3sg.poss student place-poss-dat move-pst The student whom the proctor quietly warned during the exam took his/her seat.
Scrambling Hasan’ın [makale-yi yaz-dığ-ın-ı /*yaz-dığ-ın-ı makale-yi] bil-iyor-um. Hasan-3sg.gen article-acc write-nmlz-3sg.poss-acc know-prog-1sg I know that Hasan wrote the article.
Subject Agreement [Doktor-lar /*Doktor] bu şart-ta çalış-mak zorunda değil-ler. [doctor-pl /*doctor] this condition-loc work-nmlz obliged neg-3pl [Doctors/*Doctor] do not have to work under these conditions.
Suspended Affixation Akşam kız-lar-la parti-ye [git-ti-k /*git] ve çok eğlen-di-k. evening girl-pl-com party-dat [go-pst-1pl /*go] and very have_fun-pst-1pl In the evening, we went to a party with the girls and had a lot of fun.
Table 1: Glossed minimal pairs for each phenomenon in TurBLiMP. The differences are underlined.
4.1 Phenomena
We consider 16 different grammatical phenomena, some of which are cross-lingually present in other benchmarks, alongside a few language-specific ones such as suspended affixation (see Table 1 for a complete overview with examples).
Anaphor Agreement
The anaphoric reflexive pronoun kendi agrees with its referent through number and person inflections. Unacceptable sentences in this category feature inflected forms of kendi with incorrect agreement.
Argument Structure (Transitive)
Turkish has a nominative-accusative case marking system where the direct object of a sentence is marked by the accusative case. However, a special subset of verbs assigns lexical case to their objects, deviating from structural case assignment. Unacceptable sentences feature objects with incorrect case endings, such as dative.
Argument Structure (Ditransitive)
The prototypical Turkish ditransitive construction applies a dative case marker to the indirect object. However, verbs assigning lexical case can deviate from the general trend. Here too, unacceptable sentences feature objects with incorrect case endings.
Binding
Principle B in Binding Theory (Chomsky, 1981) asserts that pronouns should be free in their binding domain, implying that pronouns should not refer to another entity in the same immediate clause. Unacceptable sentences are created by swapping an anaphora coreferring with the subject with a pronoun of similar features.
Determiners
While determiners are largely optional in Turkish, the indefinite article bir is sometimes required. When a direct object occurs immediately before the verb, its accusative case ending can be omitted. If such an object is modified by a relative clause, the indefinite article must precede the noun head (Arslan-Kechriotis, 2009). Unacceptable sentences in this phenomenon omit the obligatory determiner.
Ellipsis
This phenomenon deals with a specific type of ellipsis called backward gapping. For coordinated clauses in Turkish, it is possible to omit the verb in the first clause, leading to a gap which is resolved by the verb in the second clause. Turkish only licenses this if both clauses maintain parallel word order (Bozşahin, 2000). Acceptable sentences show the same subject-object order across clauses while unacceptable ones alternate their order.
Irregular Forms
The aorist is an aspect/ mood marker with three allomorphs -r, -Ir (high vowel harmony), and -Ar (non-high vowel harmony). While monosyllabic verbs take -Ar, a specific subset of irregular verbs take -Ir (Nakipoğlu et al., 2023). Unacceptable sentences feature an incorrect -Ar form.
Island Effects
We focus on a specific type of island constraint in which complex noun phrases are modified by a relative clause containing a wh-phrase. The occurrence of the wh-phrase is only permitted if the wh-phrase is not an adjunct (Çakır, 2016). Acceptable sentences contain argument wh-phrases like who or what, while unacceptable ones contain wh-adjuncts such as how or why.
Nominalization
Turkish extensively uses a derivational process called nominalization, where verbal bases take suffixes (like -DIK, -mA, and others) to form noun phrases. A category of Turkish verbs only selects complement clauses with -DIK, while others only allow -mA (Kornfilt, 2003b). Correspondingly, minimal pairs contain verbs with the correct and incorrect nominalization suffixes.
NPI Licensing
This phenomenon deals with Turkish negative polarity items such as hiç, kimse, hiçbir, hiçbir şey, and asla. NPIs occur in contexts where the predicate is negated. Acceptable sentences either omit the NPI or use placeholder indefinite pronouns, while unacceptable ones feature an NPI with a predicate that is not negated.
Passives
Turkish licenses the passivization of intransitive verbs via passive suffixes, creating impersonal (vs. personal) passives. While personal passives permit optional by-phrases to express agents, impersonal passives prohibit them (Özsoy, 2009). Thus, acceptable sentences omit by-phrases, while unacceptable ones include them.
Quantifiers
Turkish quantifiers such as her and çoğu can only occur with accusative-marked nouns (Enç, 1991). All minimal pairs for this phenomenon feature direct objects without accusative marking. Unacceptable sentences include a quantifier before the bare noun while acceptable sentences omit it.
Relative Clauses
Turkish uses participle suffixes -DIK and -An to form object and subject relative clauses (Göksel and Kerslake, 2005). -DIK clauses feature genitive-possessive agreement. The subject takes genitive case and the verb carries possessive agreement. In subject relative clauses with -An, only the object (if present) is case-marked. Minimal pairs target an argument preceding the nominalized verb. Acceptability depends on whether this noun is inflected with a genitive or non-genitive case ending.
Scrambling
Turkish shows word order flexibility and allows postverbal scrambling. This means that constituents can appear after the verb in certain contexts. However, local postverbal scrambling from an embedded clause is prohibited (Kornfilt, 2003a). Acceptable sentences position the object before the embedded verb while unacceptable sentences feature them in the opposite order.
Subject Agreement
Turkish realizes subject-verb agreement via person/number suffixes. Gender agreement is absent. A notable feature is third-person syncretism. The same verb inflection can indicate either a third-person singular or plural subject. However, a plural-inflected verb cannot co-occur with a singular subject. Unacceptable sentences either involve singular subjects with plural verbs or pronoun mismatches with first/second-person agreement.
Suspended Affixation
Suspended affixation refers to a phenomenon where a shared suffix applies to all conjuncts in a coordinated structure, rather than being repeated. Turkish does not allow suspended affixation for predicates inflected only with the past tense suffix -DI (Serova, 2019). Minimal pairs feature two coordinated past-tense clauses. Acceptable sentences inflect both verbs, while unacceptable ones omit inflection on the first.
4.2 Benchmark Creation
In the creation of TurBLiMP, we opted for the more labor-intensive process of manually crafting sentences. 10 initial samples per each phenomenon were created entirely manually to establish clear guidelines. This first step ensured that each pair differed only minimally while accurately capturing the targeted grammatical contrasts.
Semi-automatic augmentations
To enhance lexical diversity, we then adopted a semi-automated workflow in which a masked Turkish LM, BERTurk (Schweter, 2020) is used to suggest lexical replacements at random positions of each manually created sentence. We verified and adjusted each replacement manually to ensure acceptability. This process yielded 100 samples per phenomenon. In a final fully-automated augmentation step, BERTurk was used to generate a list of contextually appropriate words for replacement (e.g. woman or boy for girl). We use the Turkish morphology pipeline by Akın and Akın (2007) to inflect them with the same morphological features. At the end of this process, our 100 manually validated pairs increase to 1000 pairs per phenomenon. Our three-fold approach balanced scalability with linguistic precision, resulting in a robust benchmark for evaluating Turkish LMs.
4.3 Experimental Paradigms
We further assess the robustness of LMs’ syntactic abilities by focusing on two salient properties of Turkish: (i) word order flexibility and (ii) subordination through morphological processes, both discussed in Section 3. Word order variations provide a useful framework for testing the effect of word order biases on syntactic competence, extending the types of variations covered by the existing minimal pair benchmarks Linzen et al. (2016); Mueller et al. (2020). Subordination is a particularly interesting case to study the interplay between syntactic competence and morphological generalization, broadening the scope of current word-level evaluations Ismayilzada et al. (2025).
We generate word order and subordinating variations for two of the TurBLiMP phenomena (Transitive and Ditransitive Argument Structure) chosen for their flexibility for manipulation. We derive all 6 subject/verb/object orders and 4 different subordination structures for each minimal pair. Complete examples of experimental paradigms and details about how they were created are provided in Appendix A. The experimental paradigms add a total of 2,000 minimal pairs to the 16,000 pairs forming the base TurBLiMP, and considerably extend our benchmark’s utility for investigating controlled linguistic variations.
5 Human Acceptability Judgments
To validate our benchmark, we collected acceptability judgments from 30 native Turkish speakers using a 7-point Likert scale (1: completely unacceptable, 7: completely acceptable). While previous BLiMP variants rely on forced-choice tasks for data validation, BLiMP-NL (Suijkerbuijk et al., 2025) collects Likert scale responses to capture the gradient nature of acceptability judgments. We followed their approach to provide a benchmark that allows for fine-grained evaluation of model-human alignment. Our participant pool was mixed, comprising 17 linguistics students and 13 non-linguists. The study was carried out via an anonymous online survey. Appendix B includes a screenshot of survey instructions. Each participant rated 216 sentences spanning 16 linguistic phenomena as well as 20 experimental paradigms. 3 acceptable and 3 unacceptable sentences were included for each grammatical category, and the acceptability conditions were flipped between the two survey versions.
Figure 1: Mean acceptability judgments for 16 TurBLiMP phenomena. Likert scale ratings are transformed to z-scores. Error bars show standard errors of the mean.
Figure 1 reports average acceptability judgments for each phenomenon. Additional participant rating statistics are provided in Appendix C. The responses are first normalized by transforming Likert scores to z-scores. Overall, participants made clear distinctions between acceptable and unacceptable sentences. Some phenomena such as Island Effects, Passives, and Nominalization were less discriminable than others.
Phenomenon Goldfish 5MB Goldfish 10MB Goldfish 100MB Goldfish 1000MB BERTurk cosmosGPT Gemma 3 Qwen 2.5 Llama 3.1 Aya Expanse Gemma 2 EuroLLM Gemma 3 Anaphor Agreement 42.0 44.4 69.9 90.8 97.7 93.0 92.1 83.2 89.5 88.5 89.8 94.1 93.2 Argument Str. Tran. 56.1 50.6 87.9 98.3 99.1 99.3 96.5 82.5 91.3 92.8 92.2 97.6 99.1 Argument Str. Ditr. 65.3 53.2 82.5 92.3 96.1 98.0 96.6 91.7 89.6 94.6 96.8 96.7 97.6 Binding 21.1 25.8 63.3 92.3 99.0 99.2 96.5 91.0 97.1 95.6 97.9 98.6 98.2 Determiners 18.1 25.9 72.5 94.3 99.3 94.2 87.9 75.4 80.1 86.7 91.7 93.3 96.1 Ellipsis 33.0 30.2 68.0 14.9 87.5 40.2 60.5 43.6 63.8 62.7 57.5 73.2 70.4 Irregular Forms 100.0 100.0 100.0 100.0 99.6 100.0 100.0 100.0 100.0 99.4 99.9 100.0 100.0 Island Effects 93.7 92.8 81.0 78.0 51.2 89.8 75.4 86.0 78.3 72.5 75.8 79.3 75.5 Nominalization 66.4 73.9 90.8 93.3 97.4 97.0 94.9 94.0 95.3 92.4 95.9 95.2 96.6 NPI Licensing 93.4 94.3 91.9 98.1 95.0 98.5 97.7 94.0 95.4 96.9 97.1 96.9 98.1 Passives 100.0 100.0 99.9 99.9 81.3 99.4 100.0 100.0 99.6 98.8 99.5 99.7 99.9 Quantifiers 99.0 99.0 99.0 98.9 98.4 99.0 99.0 98.4 98.5 98.0 98.7 98.9 99.0 Relative Clauses 48.3 49.4 76.7 82.0 98.5 93.0 82.4 71.2 80.3 80.6 80.7 83.9 81.5 Scrambling 74.1 86.6 99.6 99.9 100.0 99.7 99.0 99.9 98.2 98.8 100.0 100.0 100.0 Subject Agreement 44.4 41.5 84.1 94.8 98.8 97.5 89.1 82.8 84.8 91.7 90.6 93.7 92.7 Suspended Affixation 57.2 64.8 93.8 98.3 100.0 99.6 99.8 97.5 98.2 98.9 99.5 99.7 100.0 Model Average 63.3 64.5 85.1 89.1 93.7 93.6 91.7 87.0 90.0 90.6 91.5 93.8 93.6 Human Correlation -0.30 -0.30 0.01 0.16 0.65 0.25 0.16 -0.07 -0.01 0.25 0.09 0.17 0.17 Parameter Count 39M 39M 125M 125M 185M 774M 4B 7B 8B 8B 9B 9B 12B Training Text Monolingual Multilingual
Table 2: Accuracy scores of each model across the linguistic phenomena in TurBLiMP. The red-green color gradient indicates performance, ranging from low to high. Significant Pearson correlations to the human judgments ($p<0.05$) are indicated in boldface.
6 Experimental Setup
Monolingual models
We employed the Goldfish series Chang et al. (2024), a series of causal LMs with fixed architecture trained on varying training data sizes (5MB, 10MB, 100MB, and 1000MB). Another monolingual model we used is BERTurk (Schweter, 2020), a 185M-parameter Turkish masked LM. With a vocabulary size of 128k, it is the only masked LM in our set of monolingual models. The largest monolingual model that we test is cosmosGPT (Kesgin et al., 2024), a 774M-parameter GPT-2-based model pretrained on Turkish web corpora and books.
Multilingual models
The evaluated multilingual models include Qwen 2.5 7B (Yang et al., 2025), Llama 3.1 8B (Meta, 2024), Aya Expanse 8B (Dang et al., 2024), Gemma 2 7B (Team et al., 2024), Gemma 3 4B and 12B (Team et al., 2025), as well as EuroLLM 9B (Martins et al., 2024). For a balanced comparison between the various models, we employed comparable parameter sizes ranging from 4B to 12B. Notably, Aya Expanse is the only instruction-tuned variant in our set of multilingual models, supporting 23 languages including Turkish. The Gemma series also boast multilinguality with Gemma 3 providing support for over 140 languages. EuroLLM prioritizes the coverage of European languages alongside a few others including Turkish.
As our evaluation metric for model performance, we computed entire-sequence log probabilities for acceptable and unacceptable sentences in each pair using the minicons library Misra (2022); Kauf and Ivanova (2023). Accuracy scores reflect the proportion of pairs where the model assigned a higher probability to the acceptable sentence. We also report Pearson’s correlation between human and model evaluations, calculated from the difference between average scores of acceptable and unacceptable sentences.
7 Results
Model performances across linguistic phenomena are summarized in Table 2. The results reveal that, more often than not, models were able to rate the acceptable sentence higher than its unacceptable counterpart. Some particular phenomena pose challenges for all the models. Ellipsis proved particularly difficult, with scores ranging from 14.9 to 87.5. Other challenging phenomena include Island Effects, Relative Clauses, and Determiners.
Figure 2: Correlation between the BERTurk model and human acceptability judgments across phenomena. (Pearson’s $r=0.65$, $p=0.007$) Each data point corresponds to the average difference per phenomenon.
Island Effects, Determiners, and Ellipsis also happen to be some of the phenomena with the lowest mean rating difference in acceptability judgments collected from native speakers as seen in Figure 1. We should note that participants preserved a clear acceptability contrast with these phenomena as well. In the case of Ellipsis, considerably low model performances are not consistent with the collected judgments. Though Ellipsis and Scrambling both manipulate word order, models handle Scrambling well. Thus, Ellipsis scores cannot be attributed to general order-manipulation difficulty.
We see that the monolingual models BERTurk and cosmosGPT tend to outperform their multilingual counterparts. Their performance is comparable to the best multilingual models EuroLLM and Gemma 3 12B. BERTurk is the only model that shows a strong cross-phenomenon correlation with human acceptability ratings, as illustrated in Figure 2. This is worth noting given that BERTurk is the only masked language model that we have tested. None of the other models had a statistically significant correlation in either direction.
Multilingual models generally show better performance with increasing model sizes, but exceptions exist. Gemma 3 4B outperforms Gemma 2 8B, and EuroLLM 9B slightly surpasses Gemma 3 12B. The superior performance of EuroLLM 9B over the same-sized Gemma 2 9B may stem from better distribution of training data across languages.
Finally, the Goldfish model series reveals the effect of training data size on performance. Models with larger training data sizes typically achieve better performance, though some counter-intuitive patterns emerge near random-chance levels. While more data generally improves learning, this pattern does not hold when acceptable sentences are consistently shorter than unacceptable ones.
7.1 Effect of Word Order
Our word order paradigm results for the best monolingual (BERTurk) and multilingual (EuroLLM) models are illustrated in Table 3. By manipulating minimal pairs for the Transitive and Ditransitive Argument Structure phenomena, we examine how different word orders affect performance.
Although SOV is the canonical word order in Turkish, Slobin and Bever (1982) found that 52% of utterances in their spontaneous adult speech corpus deviate from this order. Similarly, Türk et al. (2022) reported that only 59.5% of sentences in the BOUN Universal Dependencies Treebank follow SOV. Notably, they identified two different word orders as the second most frequent, highlighting how Turkish word order patterns can vary largely between spoken and written language. Both studies, however, agree that VOS is the least attested.
Native-speaker acceptability judgments reflect that SOV had the highest mean rating difference for both transitive and ditransitive sentences, in line with spoken and written corpus frequencies. The second-highest mean acceptability rating difference for transitive paradigms was the SVO word order, while it was OVS for the ditransitive ones. These are also the second-most-frequent word orders reported by Slobin and Bever (1982) and Türk et al. (2022) respectively. In transitive sentence ratings, VOS is not found to be the most challenging word order. This suggests that a rare word order does not inherently hinder people’s ability to identify acceptable sentences. Speakers seem to tolerate non-canonical word orders more readily in transitives than in ditransitives. One interpretation may be that case differences are easier to spot in transitive sentences due to fewer arguments.
Data Metric SOV SVO OSV OVS VSO VOS Human Slobin and Bever Freq. 48.0 25.0 8.0 13.0 6.0 0.0 BOUN Treebank Freq. 59.5 5.9 4.5 22.4 6.8 0.9 Arg. Str. Tran. $\Delta$ 1.98 1.92 1.81 1.70 1.68 1.75 Arg. Str. Ditr. $\Delta$ 1.81 1.56 1.47 1.62 1.45 1.29 EuroLLM Arg. Str. Tran. $\Delta$ 5.24 3.40 2.77 4.39 2.42 2.83 Arg. Str. Ditr. $\Delta$ 7.83 6.13 5.66 7.11 5.30 4.04 BERTurk Arg. Str. Tran. $\Delta$ 13.41 10.24 11.92 13.44 7.87 10.16 Arg. Str. Ditr. $\Delta$ 15.33 8.94 11.21 14.18 7.86 4.74
Table 3: Word order performance comparison between human judgments and best models. The white-blue gradient represents mean acceptability differences (low to high) for each row, while the white-yellow gradient reflects corpus frequency.
We see the opposite trend for model evaluations with EuroLLM being particularly sensitive to non-canonical word orders in transitive sentences. BERTurk remains robust to all word orders, showing only a pronounced drop for the rare VOS paradigm in the ditransitive condition. For both transitive and ditransitive sentences, models show high mean log probability differences on OVS word orders. This suggests that model performances align more closely with word order statistics from the BOUN treebank than with those from the spoken language corpus by Slobin and Bever (1982).
7.2 Effect of Subordination
Table 4 displays human and model performance on four subordination paradigms compared to a non-subordinated baseline. In Turkish, subordinate clauses can be finite or non-finite. However, finite subordinate clauses are much less frequent than non-finite ones (Göksel and Kerslake, 2005). For non-finite subordination, we consider three different subordinating suffixes: -DIK, -(y)IncA, and -(y)ken. -DIK forms nominal subordinate clauses while the latter two form adverbial ones.
Baseline Finite -DIK -(y)IncA -(y)ken Human Tran. $\Delta$ 1.97 1.60 1.95 0.98 0.92 Ditr. $\Delta$ 2.00 1.27 1.59 1.38 1.08 EuroLLM Tran. $\Delta$ 5.34 2.64 4.20 2.97 2.39 Ditr. $\Delta$ 6.83 4.83 5.85 5.31 4.73 BERTurk Tran. $\Delta$ 12.90 8.47 12.10 9.70 10.81 Ditr. $\Delta$ 14.14 9.45 13.43 12.35 11.74
Table 4: Subordination performance comparison between human judgments and best models.
The acceptability judgment task appears to be easier in non-finite -DIK subordinates than in finite ones, consistent with finite clauses’ lower frequency. While -DIK’s mean difference nearly matches the baseline in transitives, it shows a decline for ditransitives. Among non-finite structures, -(y)IncA and -(y)ken prove harder than -DIK, suggesting that adverbial clauses pose greater challenges. However, performance deficits may also reflect semantic incongruities from augmentation. Some verb roots may conflict with the aspectual property of the adverbial markers. Therefore, we cannot reliably claim inherent difficulty in adverbial clauses.
With human judgment patterns established, we evaluate model performance. EuroLLM’s -DIK performance shows a drop from baseline in transitive sentences. BERTurk mirrors human trends more closely, exhibiting a greater decline in ditransitives. Both models struggle more with finite subordination than -DIK, though EuroLLM shows a sharper contrast. Compared to nominal subordination, both models show smaller mean differences with adverbial clauses. Overall, we observe that models show sensitivity to different subordination structures.
8 Conclusion
TurBLiMP [^1] https://github.com/ezgibasar/turblimp provides the first comprehensive evaluation of language models’ syntactic capabilities for Turkish. We find that larger model sizes generally correlate with higher accuracy, with some exceptions. Considerably smaller monolingual language models often outperform their larger multilingual counterparts and perform on par with the best multilingual models. This finding corroborates patterns attested in other syntactic benchmarks Taktasheva et al. (2024); Jumelet et al. (2025). The strong performance of monolingual models highlights the importance of language-specific training for reliable models. Cases where smaller models outperformed larger ones also suggest that scaling alone cannot explain model behavior as far as linguistic evaluations are concerned.
The persistent challenges in phenomena like Ellipsis show that models of all sizes and architectures can struggle with some linguistic phenomena. The discrepancy between model behavior and human judgments for this phenomenon indicates that even the best-performing LLMs may fail to fully capture human linguistic intuition.
TurBLiMP also introduces experimental paradigms to test model robustness to specific linguistic parameters, namely word order and subordination. Results on these paradigms reveal subtle sensitivities in high-performing models that standard evaluations would miss, indicating that even models excelling on general minimal pair tasks can exhibit brittleness with the introduction of non-canonical word orders or subordination. While some performance patterns align with human judgments, we observe both variation across models and cases of divergence from human judgments. Furthermore, the human judgments themselves offer valuable insights into native speaker patterns across different subordination structures and word orders, making TurBLiMP a valuable starting point for future research.
In sum, TurBLiMP provides a valuable resource to assess various linguistic phenomena in a controlled fashion, many of which are not represented in prior syntactic benchmarks. We hope our work will facilitate linguistically informed model developments and contribute to a better understanding of how language models handle linguistic structures across typologically different languages.
Limitations
While TurBLiMP offers a comprehensive evaluation of key linguistic phenomena in Turkish, there are also several limitations to acknowledge. In this paper, we evaluated minimal pair acceptability using sequence log probabilities for each sentence. However, our approach represents only one of several valid methods for assessing language models on acceptability benchmarks. For example, Song et al. (2025) show that prompting for ‘meta-linguistic’ grammaticality judgments can result in better performance than comparing string probabilities directly. However, they do show that this ‘introspective’ approach has its limitations, and the optimal way of evaluating linguistic ability in LMs remains an open debate (Hu et al., 2024).
Warstadt et al. (2020) define paradigms as minimal pair types and phenomena as broader linguistic categories. Unlike other BLiMP benchmarks that include multiple paradigms under each phenomenon, TurBLiMP currently includes only one paradigm per phenomenon (with the exception of Argument Structure). Splitting some phenomena into multiple paradigms could enable more granular assessment. For a broader coverage of linguistic structures in Turkish, future work could also incorporate additional phenomena into the benchmark.
Our evaluation did not systematically test all available model sizes for each model, which may limit the generalizability of our findings as far as model size is concerned. Testing a wider range of model sizes would strengthen our insights.
Some of the models we tested had both base and instruction-tuned variants. We excluded instruction-tuned variants as the English-based tuning was unlikely to improve performance on Turkish linguistic tasks. However, this design choice means we cannot speak to potential transfer effects from instruction tuning.
Our word order experiments focused exclusively on sentences with explicit subjects, omitting pro-drop constructions. Given Türk et al. (2022)’s finding that subjectless sentences exceed canonical SOV order frequency in the BOUN treebank, future work should investigate these prevalent but untested configurations.
Appendix A Experimental Paradigm Details
Our experimental paradigms only target minimal pairs belonging to Argument Structure Transitive and Argument Structure Ditransitive phenomena. Below is an example of the transitive baseline.
[…]
Bu [şarkı-ya /*şarkı-yı] bayıl-ıyor-um.
this [song-DAT /*song-ACC] love-PROG-1SG
[…]
‘I love this song.’
For word order, we took the 100 manually crafted minimal pairs from the Transitive and Ditransitive Argument Structure phenomena and, for each phenomenon, we generated 600 pairs for all six possible subject-verb-object permutations. While the original sentences predominantly followed the default Turkish SOV order, the subjects were often omitted and left implicit. We explicitly reintroduced subjects in a dedicated SOV variant and derived the remaining five orders from this augmented set.
[…]
Data Metric SOV SVO OSV OVS VSO VOS EuroLLM Arg. Str. Tran. Acc. 97.0 91.0 81.0 89.0 82.0 84.0 Arg. Str. Ditr. Acc. 98.0 97.0 92.0 94.0 95.0 86.0 BERTurk Arg. Str. Tran. Acc. 99.0 98.0 99.0 100.0 96.0 96.0 Arg. Str. Ditr. Acc. 98.0 96.0 96.0 98.0 95.0 74.0
Table 5: Accuracy scores for word order paradigms.
For subordination, we augmented each paradigm by creating subordinate clauses with three different subordinating suffixes (-(y)IncA, -(y)ken, -DIK) and included a finite subordination paradigm which preserves the original verb inflection. -DIK forms nominal subordinate clauses, while -(y)IncA and -(y)ken form adverbial subordinate clauses. -DIK nominalization carries agreement and case suffixes, and -(y)ken attaches to verb stems marked for aspect.
[…]
Finite
Bu [şarkıya /*şarkıyı] bayılıyorum sanıyor.
[…]
Bu [şarkıya /*şarkıyı] bayıldığımı sanıyor.
[…]
Bu [şarkıya /*şarkıyı] bayılınca gitti.
[…]
Bu [şarkıya /*şarkıyı] bayılırken gitti.
$\sim$(S)he left while I was loving this song.’
These strategies show varying degrees of morphological complexity, with finite subordination using 2 morphemes, -DIK using 3, -(y)IncA using 1, and -(y)ken using 2. The augmentation procedure yields 400 transitive and 400 ditransitive subordination minimal pairs.
Baseline Finite -DIK -(y)IncA -(y)ken EuroLLM Tran. Acc. 97.6 82.0 89.0 89.0 77.0 Ditr. Acc. 96.7 89.0 97.0 94.0 88.0 BERTurk Tran. Acc. 99.1 97.0 98.0 98.0 97.0 Ditr. Acc. 96.1 91.0 95.0 96.0 96.0
Table 6: Accuracy scores for subordination paradigms.
Appendix B Acceptability Judgment Collection
Figure 3 provides a screenshot of the survey carried out on the Qualtrics platform. All participants gave their informed consent before starting the survey and agreed that their anonymous responses can be made publicly available. Students received extra credit in exchange for their participation.
Figure 3: Informed consent form and instructions.
Appendix C Acceptability Judgment Statistics
Category Proportion
Transitive SOV 30/30
Transitive SVO 30/30
Transitive OSV 30/30
Transitive OVS 29/30
Transitive VSO 30/30
Transitive VOS 30/30
Ditransitive SOV 30/30
Ditransitive SVO 30/30
Ditransitive OSV 28/30
Ditransitive OVS 28/30
Ditransitive VSO 30/30
Ditransitive VOS 29/30
Table 7: Proportion of participants with higher average acceptability ratings for acceptable versus unacceptable sentences per word order paradigm.
Category Proportion
Transitive Baseline 30/30
Transitive Finite 30/30
Transitive -DIK 30/30
Transitive -(y)IncA 24/30
Transitive -(y)ken 28/30
Ditransitive Baseline 29/30
Ditransitive Finite 29/30
Ditransitive -DIK 30/30
Ditransitive -(y)IncA 29/30
Ditransitive -(y)ken 28/30
Table 8: Proportion of participants with higher average acceptability ratings for acceptable versus unacceptable sentences per subordination paradigm.
Category Proportion
Anaphor Agreement 30/30
Argument Structure Tran. 30/30
Argument Structure Ditr. 29/30
Binding 30/30
Determiners 28/30
Ellipsis 27/30
Irregular Forms 30/30
Island Effects 26/30
Nominalization 30/30
NPI Licensing 30/30
Passives 29/30
Quantifiers 30/30
Relative Clauses 30/30
Scrambling 30/30
Subject Agreement 30/30
Suspended Affixation 30/30
Table 9: Proportion of participants with higher average acceptability ratings for acceptable versus unacceptable sentences per phenomenon.